# IQ-Aware Precoding for Atomic MIMO Receivers

### v1

@@ VI-C Validation of Strong Reference Approximation
Figure 6: The mutual information between ${\mathbf{y}}$ and ${\mathbf{x}}$ against RSNR for the magnitude-detection model in (5) and real-part-detection model in (8).
@@
VI-D Achievable Rate of Proposed Precoding Schemes
@@
The curves in Fig. 7 compare different transmitter precoding approaches. The achievable rate of atomic MIMO receivers in (13) is adopted for performance evaluation. The benchmarks include: 1) the optimal IQ-aware digital precoding; 2)-3) the proposed IQ-aware hybrid precoding for the FC- and SC-PS structures; 4) the classic digital precoding, which assigns the complex precoder ${\mathbf{F}}$ with the principal singular vectors of the complex channel $\tilde{{\mathbf{H}}}$ associated with water-filling principle; 5)-6) the classic hybrid precoding algorithms, PE-AltMin for the FC structure and SDR-AltMin algorithm for the SC structure [34]. For both sub-figures, the number of data streams, RF chains, and transmit antennas are set to $N_{s}=3$, $N_{\rm RF}=12$, and $N_{t}=48$. In Fig. 3 (a), the number of receive antennas is $N_{r}=72$ and the receive SNR grows from -10 dB to 5dB, while in Fig. 3 (b), the receive SNR is fixed as 0 dB and $N_{r}$ increases from 32 to 96. We run 1000 Monte-Carlo trails to plot each curve.
We can observe that all proposed IQ-aware precoding schemes outperform their traditional counterparts for all considered SNRs and the number of receive antennas. For example, when ${\rm SNR}=0$ dB and $N_{r}=72$, the gap in achievable rate is approximately 2 bit/s/Hz for the FC hybrid precoding architecture and around 1 bit/s/Hz for the SC hybrid precoding architecture, demonstrating the effectiveness of the proposed design. One can also observe that with the increase in SNR, the rate achieved by IQ-aware hybrid precoding (FC) gradually approaches that of the optimal IQ-aware digital precoding. This comes from the fact that the power allocation to the orthogonal columns of the digital precoder $\bar{{\mathbf{F}}}$ tend to be uniform at high SNRs, making the orthogonal-column assumption applied on $\bar{{\mathbf{D}}}$ more precise. Consequently, we can conclude that an analog FC-PS network associated with a low-dimensional IQ-aware digital precoder is able to accurately approximate a high–dimensional IQ-aware digital precoder.
@@ VI-E Convergence Verification
Figure 8: The objective functions, (51) and (71), w.r.t the number of iterations.
The curves in Fig. 8 demonstrate the convergence of Algorithm 1 and Algorithm 2. The value of objective functions, $\|\widetilde{{\mathbf{F}}}\widetilde{{\mathbf{D}}}^{T}-\bar{{\mathbf{A}}}\|_{F}^{2}$ and $\|\bar{{\mathbf{F}}}-\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\|_{F}^{2}$, in each iteration are plotted in Fig. 8(a) and (b), respectively. The simulation parameters are as follows: $N_{s}=2$, $N_{t}=32$, $N_{r}=64$, and ${\rm SNR}=0$ dB. It is clear that our proposed algorithms can monotonically decrease the objective functions until convergence. The number of iterations required for convergence is around 50 for Algorithm 1 and is around 10 for Algorithm 2. The larger iteration number of Algorithm 1 comes from the updating of auxiliary variables. Moreover, since the FC structure has much more PSs, the objective function of Algorithm 1 can converge to 0 when $N_{\rm RF}=8$, while the converged value of Algorithm 2 is larger than 0 when $N_{\rm RF}=16$.
VII Conclusion
In this paper, we explored transmitter precoding designs for approaching the capacity of Rydberg atomic receivers in MIMO systems. We proposed a strong reference approximation to linearize the magnitude-detection model, effectively converting it into a real-part detector. Building on this, we established the limit of ergodic capacity as $\min\{N_{r}/2,N_{t}\}\log\{\rm SNR\}+\mathcal{O}(1)$, such that a DoF-lossless transmission is enabled by atomic receivers when $N_{r}\geq 2N_{t}$. Then, we proposed an IQ-aware fully digital precoding to achieve the instantaneous capacity of atomic receivers, which offers more flexible manipulation on the IQ baseband symbols. To facilitate the implementation of analog RF circuits, we further delved into the IQ-aware hybrid precoding algorithms for both FC and SC architectures. Our simulation results confirmed that the proposed IQ-aware precoding schemes significantly outperform traditional precoding methods.
This work presents a significant stride towards the transmitter-side signal processing for enhancing atomic MIMO receivers. We conclude by outlining potential avenues for future research. Firstly, this study was limited to point-to-point MIMO systems. However, in the context of multi-user communications where each user utilizes an atomic receiver, it becomes crucial to consider inter-user interference. In such a scenario, the proposed precoding schemes may not be optimal. Secondly, our analyses and designs were built on the strong reference approximation. In case the LO is absent or the strong reference approximation is invalid, maximizing throughput for the nonlinear transition model becomes necessary. This is still an open problem that requires further investigation.
-A Proof of Proposition 1
@@ VII Conclusion
-B Proof of Proposition 2
When ${\rm SNR}\rightarrow+\infty$, the ergodic capacity tends to $\frac{1}{2}\mathsf{E}\left[\log\det\left(\frac{\rm SNR}{N_{t}}\bar{{\mathbf{H}}}\bar{{\mathbf{H}}}^{T}\right)\right]=\frac{1}{2}\mathsf{E}\left[\log\det\left(\frac{\rm SNR}{N_{t}}\bar{{\mathbf{H}}}^{T}\bar{{\mathbf{H}}}\right)\right]$. We denote $m=\min\{N_{r}/2,N_{t}\}$ and $M=\max\{N_{r}/2,N_{t}\}$, and further define
${\mathbf{W}}=\left\{\begin{array}[]{c}\bar{{\mathbf{H}}}\bar{{\mathbf{H}}}^{T}\quad\quad{\rm if}\>\>N_{r}\leq 2N_{t}\\ \bar{{\mathbf{H}}}^{T}\bar{{\mathbf{H}}}\quad\quad{\rm if}\>\>N_{r}>2N_{t}\end{array}\right..$ (100)
Since the coefficients of $\bar{{\mathbf{H}}}\in\mathbb{R}^{{N_{r}}\times{2N_{t}}}$ follow the i.i.d Gaussian distribution $\mathcal{N}(0,\frac{1}{2})$, ${\mathbf{W}}\in\mathbb{R}^{2m\times 2m}$ has a Wishart distribution with $2M$ degrees of freedom: ${\mathbf{W}}\sim W_{2m}\left({\mathbf{V}},2M\right)$, where ${\mathbf{V}}=\frac{1}{2}{\mathbf{I}}_{2m}$. Substituting ${\mathbf{W}}$ into the log-expectation expression of Wishart distribution [38], we arrive at
$\frac{1}{2}\mathsf{E}\left[\log\det\left(\frac{\rm SNR}{N_{t}}{\mathbf{W}}\right)\right]=m\log\frac{\rm SNR}{N_{t}}+\frac{1}{2}\mathsf{E}\left[\log\det\left({\mathbf{W}}\right)\right]$
$=m\log\frac{\rm SNR}{N_{t}}+\frac{1}{2}\left(\log\left|\frac{1}{2}{\mathbf{I}}_{2m}\right|+2m\log(2)+\psi_{2m}(M)\right)$
$=m\log\frac{\rm SNR}{N_{t}}+\frac{1}{2}\psi_{2m}(M),$ (101)
where $\psi_{2m}(M)$ denotes the multivariate digamma function (the logarithmic derivative of the multivariate gamma function). Since $\frac{1}{2}\psi_{2m}(M)$ is a constant independent of SNR, the ergodic capacity is asymptotically $\min\{N_{r}/2,N_{t}\}\log({\rm SNR}/{N_{t}})+\mathcal{O}(1)$ at high SNRs. This completes the proof.
-C Proof of Lemma 1

### v2

@@ VI-C Validation of Strong Reference Approximation
Figure 6: Mutual information between ${\mathbf{y}}$ and ${\mathbf{x}}$ against RSNR for the magnitude-detection model in (5) and real-part-detection model in (8).
@@
VI-D Achievable Rates of Proposed IQ-aware Precoding
@@
The curves in Fig. 7 compare different transmitter precoding approaches. The achievable rate of atomic MIMO receivers in (12) is adopted for performance evaluation. The benchmarks include: 1) the optimal IQ-aware digital precoding; 2)-3) the proposed IQ-aware hybrid precoding for the FC- and SC-PS structures; 4) the classic digital precoding, which assigns the complex precoder ${\mathbf{F}}$ with the principal singular vectors of the complex channel $\widetilde{{\mathbf{H}}}$ associated with water-filling principle; 5)-6) the classic hybrid precoding algorithms, PE-AltMin for the FC structure and SDR-AltMin algorithm for the SC structure [39]. For both sub-figures, the number of data streams, RF chains, and transmit antennas are set to $N_{s}=3$, $N_{\rm RF}=12$, and $N_{t}=48$. In Fig. 2 (a), the number of receive antennas is $N_{r}=12$ and the receive SNR grows from -5 dB to 10 dB, while in Fig. 2 (b), the receive SNR is fixed as 0 dB and $N_{r}$ increases from 8 to 32. We run 1000 Monte-Carlo trails to plot each curve.
We can observe that all proposed IQ-aware precoding schemes outperform their traditional counterparts for all considered SNRs and the number of receive antennas. For example, when ${\rm SNR}=0$ dB and $N_{r}=12$, the gap in achievable rate is approximately 3 bit/s/Hz for the FC hybrid precoding architecture and around 2 bit/s/Hz for the SC hybrid precoding architecture, demonstrating the effectiveness of the proposed design. One can also observe that with the increase in SNR, the rate achieved by IQ-aware hybrid precoding (FC) gradually approaches that of the optimal IQ-aware digital precoding. This comes from the fact that the power allocation to the orthogonal columns of the digital precoder $\bar{{\mathbf{F}}}$ tend to be uniform at high SNRs, making the orthogonal-column assumption applied on $\bar{{\mathbf{D}}}$ more precise. Consequently, we can conclude that an analog FC-PS network associated with a low-dimensional IQ-aware digital precoder is able to accurately approximate a high–dimensional IQ-aware digital precoder.
@@ VI-E Convergence Verification
Figure 8: Objective functions, (47) and (66), w.r.t the number of iterations.
The curves in Fig. 8 demonstrate the convergence of Algorithm 1 and Algorithm 2. The value of objective functions, $\|\widetilde{{\mathbf{F}}}\widetilde{{\mathbf{D}}}^{T}-\bar{{\mathbf{A}}}\|_{F}^{2}$ and $\|\bar{{\mathbf{F}}}-\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\|_{F}^{2}$, in each iteration are plotted in Fig. 8(a) and (b), respectively. The simulation parameters are as follows: $N_{s}=2$, $N_{t}=32$, $N_{r}=12$, and ${\rm SNR}=0$ dB. It is clear that our proposed algorithms can monotonically decrease the objective functions until convergence. The number of iterations required for convergence is around 50 for Algorithm 1 and is around 10 for Algorithm 2. The larger iteration number of Algorithm 1 comes from the updating of auxiliary variables. Moreover, since the FC structure has much more PSs, the objective function of Algorithm 1 can converge to 0 when $N_{\rm RF}=8$, while the converged value of Algorithm 2 is larger than 0 when $N_{\rm RF}=16$.
VII Conclusions
In this paper, we explored precoding design for achieving the capacity of Rydberg atomic receivers in MIMO systems. We proposed a strong-reference approximation to linearize the magnitude detection model into a real-part detection model. An IQ-aware digital precoding is accordingly designed to show that the channel capacity of atomic MIMO receiver scales as $\min(N_{r}/2,N_{t})\log(\rm SNR)+\mathcal{O}(1)$, revealing a DoF-lossless transmission when $N_{r}\geq 2N_{t}$. To facilitate the implementation of analog RF circuits, we further delved into the IQ-aware hybrid precoding algorithms for both FC and SC architectures. Simulation results confirmed that the proposed IQ-aware hybrid precoding can tightly approach the optimal IQ-aware digital precoding.
This work presents a significant stride towards the transmitter-side signal processing for enhancing atomic MIMO receivers. We conclude by outlining potential avenues for future research. Firstly, this study was limited to point-to-point MIMO systems. In the context of multi-user communications where each user utilizes an atomic receiver, it becomes crucial include inter-user interference into the IQ-aware precoding design. Secondly, our analyses and designs were built on the strong-reference approximation. In case the LO is absent or the strong-reference approximation is invalid, maximizing the mutual information of nonlinear transition model becomes necessary, which remains an open problem. Last but not the least, precoding design for the general heterodyne measurement, where the frequency of LO slightly differs from that of signal [19], is an unexplored area as well.
-A Proof of Lemma 1
@@ VII Conclusions
-B Proof of Lemma 3